Appendix B Present value of $1, PV, PV = FV (1+1 Period 10% 0.909 0.826 0.751 0.683 1% 0.990 0.980 0.971 0.961 0.951 0.942 0.(1 + i) Appendix D Present value of an annuity of $1, PVC PVA Period 1% 0.990 1.970 2.941 3.902 4.853 5.795 6.728 7.652 8.566Media Bias Inc. issued bonds 10 years ago at $1,000 per bond. These bonds had a 30-year life when issued and the annual interest payment was then 11 percent. This return was in line with the required returns by bondholders at that point in time as described below: Real rate of return 2 % Inflation premium 4 Risk premium 5 Total return 11 % Assume that 10 years later, due to good publicity, the risk premium is now 3 percent and is appropriately reflected in the required return (or yield to maturity) of the bonds. The bonds have 20 years remaining until maturity.
Answer #1

Using Appendix:
Price now=1000*11%*P/A(11%-5%+3%,20)+1000*P/F(11%-5%+3%,20)=1000*11%*P/A(9%,20)+1000*P/F(9%,20)=1000*11%*9.129+1000*0.178=1182.19

Using formula:
Price now=1000*11%/9%*(1-1/1.09^20)+1000/1.09^20=1182.570913

Using financial calculator:

I/Y=9%

N=20

PMT=11%*1000=110

FV=-1000

CPT PV=1182.570913